Calculate the YTM of a perpetuity when it has a price of $15,000 and an annual coupon payment on $900.
Answer #1

Let YTM be i

Current price=P=$15000

Annual coupon payment=C=$900

Present worth of perpetuity=C/i=900/i

We know that current price of perpetuity is equal yo Present worth of perpetuity. So,

15000=900/i

i=900/15000=6%

YTM is 6%